OpenStudy (lgbasallote):

could someone please give me an integration problem to solve? (no drawings involved :P)

OpenStudy (unklerhaukus):

\[\int\limits_0^1 e^x+2x+100\quad\text dx\]

OpenStudy (anonymous):

http://www.scribd.com/doc/33405554/Integration-Practice-Test-1 <---bc I can't do it myself ;p

OpenStudy (anonymous):

integrate x^2/4+x^2

OpenStudy (amistre64):

\[\int e^{-x^2}dx\]

OpenStudy (marco26):

\[\int\limits \sin^4(4y)\cos^6(4y) dy\]

OpenStudy (lgbasallote):

@amistre64 i think that's unintegrable? :P

OpenStudy (lgbasallote):

@UnkleRhaukus is the answer e + 100?

OpenStudy (anonymous):

if \[C _{r}=\left(\begin{matrix}10 \\ r\end{matrix}\right), then \sum_{r=1}^{10}\]\[r.C _{r}\over C _{r-1} \] is

OpenStudy (amistre64):

:) its integrable, but you have to do it in another way

OpenStudy (lgbasallote):

how did that become integration @Rohangrr o.O

OpenStudy (unklerhaukus):

@lgbasallote almost \[=e+101\]

OpenStudy (lgbasallote):

it's e + 101 - 1 right?

OpenStudy (anonymous):

@Rohangrr: I have no idea why are posting off-topic things in various threads, would you like to enlighten me?

OpenStudy (unklerhaukus):

why minus one?

OpenStudy (anonymous):

@lgbasallote and @FoolForMath i had doubt on that!! pls help!! :)

OpenStudy (lgbasallote):

[e^1 + 1^2 + 100(1)] - [e^0 + 0^2 + 100(0)] [e + 1 + 100] - [1 + 0 + 0] e + 101 - 1 e + 100

OpenStudy (lgbasallote):

@Rohangrr im asking problems myself :P dont ask me of your problems lol..@amistre64 im feeling im getting warm

OpenStudy (lalaly):

\[\int\limits{\sqrt{tanx}}dx\]

OpenStudy (amistre64):

\[\lim_{n \to\ \infty}\sum_{k=1}^{n} \frac{3}{n}(\frac{3k}{n})^{2}\] i think i notated it right

OpenStudy (lgbasallote):

lol @amistre64 i reached \(\LARGE \int \frac{ds}{e^{s^2}}\) went in circles after 2 subs

OpenStudy (lgbasallote):

im gonna solve lana's problem

OpenStudy (unklerhaukus):

oh your right @lgbasallote my mistake very good work

OpenStudy (anonymous):

For @amistre64 first problem I would use numerical analysis.

OpenStudy (lgbasallote):

hehe :DDD aww numerical analysis >.<

OpenStudy (anonymous):

For the second one, you form the integration.

OpenStudy (amistre64):

that is the integration form

OpenStudy (lgbasallote):

you drive a hard bargain @lalaly -___-

OpenStudy (amistre64):

the interval is divided up infinitely; and the function is summed up at all the right end points

OpenStudy (lalaly):

no thats not right

OpenStudy (amistre64):

Waheed, not quite

OpenStudy (lgbasallote):

could i get a hint on how to integrate \(\int \sqrt{\tan x} dx?\) @lalaly

OpenStudy (amistre64):

\[\int sin(x^2)dx\]

OpenStudy (lgbasallote):

lol is that another numerical value thingy amistre :P

OpenStudy (amistre64):

;)

OpenStudy (amistre64):

the integrations they give you to start with in textbooks is to build your confidence and to get you to believe in the processes; but there are infinitely many integrations that cannot simply be done

OpenStudy (lgbasallote):

that's a nice way to pop my fantasy bubble amistre :P

OpenStudy (amistre64):

its a hobby ;)

OpenStudy (anonymous):

2/3(cube root (tanx)+lnsecx+c answer of squar

OpenStudy (lalaly):

let u^2=tanx

OpenStudy (marco26):

algebraic substitution?

OpenStudy (lgbasallote):

i got \(\Large 2 \int \frac{u^3}{1+u^4} du\)

OpenStudy (lgbasallote):

lol crazy manipulation @lalaly :P

OpenStudy (anonymous):

@amistre can u solve the first question gave in here ....

OpenStudy (lgbasallote):

hmmm this thread made me realize...i need to learn how to use partial fractions in integration

OpenStudy (anonymous):

hey @lalay can u solve the question u gave here